OpenStudy (anonymous):

Factor 2x2 + 7x + 3. A. (2x + 1)(x + 3) B. (2x + 3)(x + 1) C. (2x + 5)(x + 1) D. The polynomial is prime.

OpenStudy (anonymous):

i think A

OpenStudy (the_fizicx99):

2x^2 + 7x + 3 2 * 3 = 6 1 * 6 = 6 1 + 6 = 7 2x^2 + x + 6x + 3 x(2x + 1) + 3(2x + 1) (2x + 1)(x + 3)

OpenStudy (anonymous):

So i was correct?

OpenStudy (the_fizicx99):

Yes.

OpenStudy (anonymous):

Factor 3x2 – 13x + 4. A. (3x – 1)(x – 4) B. (3x – 2)(x – 2) C. (3x – 4)(x – 1) D. The polynomial is prime.

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Show me your work?

OpenStudy (anonymous):

\[3x ^{2}-13x+4(3x-1)(x-4)\] =3x+12+1x+4 FOIL METHOD =13x+4

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Um not really. Take the first coefficient and multiply it by the constant in the back. \(\ 3 \times 4 \) = ?

OpenStudy (anonymous):

12

OpenStudy (the_fizicx99):

Yes, now the semi hard part is find one number that when multiplied gives you 12 but adds up to your middle term which in this case is -13

OpenStudy (the_fizicx99):

No, what number when multiplied gives you 12 but adds up to -13?

OpenStudy (anonymous):

never mind thats not correct

OpenStudy (the_fizicx99):

Hint: Use negative numbers Ex: -1 * -5 = 5 But -1 + (-5) = -1 - 5 = -6

OpenStudy (the_fizicx99):

@milbes919180 ?

OpenStudy (the_fizicx99):

You sure on that?

OpenStudy (the_fizicx99):

I have to get off. -1 * -12 = 12 -1 - 12 = -13 3x^2 - x - 12x + 4 x(3x - 1) -4(3x -1) (3x - 1)(x -4)

OpenStudy (anonymous):

Thats what i answered the first time

OpenStudy (the_fizicx99):

You said -12 and 1...

OpenStudy (anonymous):

3x2−13x+4(3x−1)(x−4) =3x+12+1x+4 FOIL METHOD =13x+4

OpenStudy (the_fizicx99):

That's wrong. It doesn't work like that.

OpenStudy (anonymous):

u coulda said right answer wrong work

OpenStudy (the_fizicx99):

How did you get that answer, if I may ask?

OpenStudy (anonymous):

do you not see the (3x-1)(x-4)

OpenStudy (the_fizicx99):

I see it. I mean how did you get that answer if you did the process incorrectly? You used the options. Even if you did use the options, you didn't verify them correctly.